OpenStudy (anonymous):

find the quotient. (49w^5-7w) / 7w explain.

Nnesha (nnesha):

on the top you have first thing common what is that numerator

OpenStudy (shinalcantara):

\[\frac{ 49w^5 -7w }{ 7w } = \frac{ 49w^5 }{ 7w } - \frac{ 7w }{ 7w }\] simplify

Nnesha (nnesha):

 let me know what is common ???

OpenStudy (anonymous):

hmm well the letters lol

Nnesha (nnesha):

and what else ???

OpenStudy (anonymous):

well 7x7 is 49

Nnesha (nnesha):

yep 7w is common

OpenStudy (anonymous):

so 7w/7?

Nnesha (nnesha):

 now fill the bracket